OpenStudy (anonymous):

The probability of event A occurring is 50%, and the probability of event B occurring is 25%. Events A and B are nonsimultaneous events. What is the probability that event A occurs but event B does not occur?

OpenStudy (ankit042):

p(a) = 0.5 p(b) = 0.25 p(not a) = 1-0.5 = 0.5 p(not b) = 1 - 0.25 = 0.75 Now p(a and not b) = p(a) * p(not b) You can get the answer now!

OpenStudy (anonymous):

so the answer is 0.375?

OpenStudy (anonymous):

Thank YOU!!!!!

OpenStudy (ankit042):

0.5 *0.75 = 0.375 np :)